OpenStudy (anonymous):

Describe the graph of the function f(x) = x3 − 18x2 + 101x − 180. Include the y-intercept, x-intercepts, and the shape of the graph. Can you give me then answer and tell me how you got it and walk me through it.

OpenStudy (welshfella):

the y intercept occurs when x = 0 so what do you think the y-intercepts is?

OpenStudy (anonymous):

how would I find the Y intercept?

OpenStudy (welshfella):

plug in x = 0 and work it out

OpenStudy (welshfella):

0^3 - 18(0)^2 + 101(0) - 180 = ?

OpenStudy (anonymous):

y=180?

OpenStudy (welshfella):

-180

OpenStudy (anonymous):

so x=0 and y=-180 how would I figure out the shape of the graph?

OpenStudy (welshfella):

well try and find the x intercepts first they occur when y = 0 so we have the equation x^3 − 18x^2 + 101x − 180 = 0

OpenStudy (anonymous):

so far I got 180=x^3 - 18^2 + 101x

OpenStudy (welshfella):

- solve for x by the rational roots theorem the roots could be factors 0f 180 like +/- 1, +/- 3, +/- 6 , +/-5,+/- 6, +/- 9 so there's a lot to try out!

OpenStudy (welshfella):

or we might be able to factor by grouping which would be a lot quicker:- x3 − 18x2 + 101x − 180

OpenStudy (anonymous):

I got 3 answers and I don't know which one is right x = 9 x = 5 x = 4

OpenStudy (welshfella):

nope cant do it that way

OpenStudy (welshfella):

how did you get those?

OpenStudy (welshfella):

- they might all be right. A cubic equation will have 3 roots.

OpenStudy (welshfella):

f(5) = 5^3 - 18(25) + 101(5) - 180 = 0 yes 5 is a root

OpenStudy (anonymous):

im so confused, so x would be 5 and y is -180

OpenStudy (welshfella):

did you get a calculator or wolfram to do it?

OpenStudy (welshfella):

no x = 5 when y = 0

OpenStudy (welshfella):

how did you get those 3 roots They are all correct I just checked them on my calc.

OpenStudy (anonymous):

I couldn't figure it out so I used a calculator

OpenStudy (welshfella):

oh i see well another way is to to try ouy values such as 3, 5 etc and once you have found one - say 5- you divide the function by x - 5 using long or synthetic division to get a quadratic which you solve to get the other 2.

OpenStudy (welshfella):

to find the shape you can differentiate and find slopes at different values of x We know it is probably a double loop because there are 3 rational roots.

OpenStudy (anonymous):

so the shape is a loop?

OpenStudy (welshfella):